OpenStudy (anonymous):

how can i solve this; -3sqrt24-sqrt24+3sqrt5+3sqrt54

OpenStudy (anonymous):

= -4\[\sqrt{24} + 3\sqrt{5} + 9\sqrt{6}\]

OpenStudy (anonymous):

sorry, I typed the worng one

OpenStudy (gw2011):

-3(square root 24)-square root 24+33(square root 5)+3square root 54) -4(square root 24)+3(square root 5)+3(square root 9*square root 6) -4(square root 4*square root 6)+3(square root 5)+3(square root 9*square root 6) (-4)(2)(square root 6)+3(square root 5)+(3)(3)(square root 6) -8(square root 6)+3(square root 5)+9(square root 6) square root 6+3(square root 5)